OpenStudy (anonymous):

Need someone to check my answer! A website recorded 1,351 visitors one month. This was 30% less than the number of visitors to the site the previous month. The following equation represents this situation, where x represents the number of visitors to the site the previous month. x − 0.3x = 1,351 How many visitors did the website record in both months combined?

OpenStudy (anonymous):

i got 1930, is that correct?

OpenStudy (anonymous):

x − 0.3x = 1351 , x = 1351/0.7 = 1930. The number of visitors in both months combined = 1351 + x = 1351 + 1930 = 3281 visitors.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

You're welcome.